Circular No. 7632

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

1999 KW_{4}

L. A. M. Benner, S. J. Ostro, J. D. Giorgini and R. F. Jurgens, Jet Propulsion Laboratory; J. L. Margot, California Institute of Technology; and M. C. Nolan, Arecibo Observatory, write: “Goldstone (8560-MHz, 3.5cm) delay-Doppler radar images on May 21–23 reveal that 1999 KW_{4} is a binary system. The images show separations up to at least 2 km between the components, whose sizes differ by a factor of at least three. Additional radar observations are planned at Goldstone and Arecibo. Follow-up optical observations are desirable.”

The following ephemeris is from the orbital elements on MPEC 2001-K23 (2 km at the object’s distance corresponding to 0′′.06–0′′.09):

2001 TT α_{2000} δ_{2000} Δ r ϵ β V
May 24.0 20^{h}46.^{m}52 −18^{o}28.^{′}3 0.040 1.029 113.^{o}8 64.^{o}1 11.9
24.5 20 18.18 −15 19.5 0.037 1.032 120.0 58.2 11.6
25.0 19 46.27 −11 23.8 0.034 1.034 126.8 51.7 11.3
25.5 19 11.52 − 6 45.0 0.033 1.036 133.8 44.9 11.0
26.0 18 35.28 − 1 39.0 0.032 1.038 139.9 38.9 10.8
26.5 17 59.30 + 3 28.4 0.033 1.040 144.2 34.7 10.7
27.0 17 25.27 + 8 11.4 0.035 1.042 145.9 33.0 10.8
27.5 16 54.36 +12 13.6 0.037 1.044 145.2 33.6 10.9
28.0 16 27.12 +15 30.5 0.040 1.046 143.0 35.7 11.2
28.5 16 03.59 +18 05.9 0.044 1.048 140.1 38.3 11.5
29.0 15 43.45 +20 06.7 0.048 1.049 137.2 41.1 11.7

NOVA AQUILAE 2001 CCD magnitude estimate by J. D. West, Mulvane, KS: May 19.18 UT,

V = 12.94 ± 0.05.

K. Hornoch, D. Hanžl and M. Lehký, Brno, Czech Republic, provide the following CCD photometry on May 20.04 UT: V = 13.27, R = 11.5, I = 10.47 (comparison star HIP 93914: V = 8.73, B − V = +0.339 ± 0.024, V − I = +0.40 ± 0.03). During May 20.029–20.086 the nova was constant in R to within ±0.01 mag.

Visual magnitude estimates provided by J. Mattei, American Association of Variable Star Observers: May 20.98 UT, 13.7 (P. Dubovsky, Podbiel, Slovakia); 21.35, 13.8 (P. Abbott, Leduc, AB, Canada).

Corrigendum. On IAUC 7628, line 1, for Instituto de Astrofísica de Canarias read Isaac Newton Group, La Palma

2001 May 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Brian G. Marsden